If three resistor are connected in series, what is the voltage drop in R1 if the value of R1= 12 ohm R2 = 20 ohm R3 = 5 ohm and the voltage source is Et = 24 volts


1
Expert's answer
2021-03-29T14:35:59-0400

The total resistance is the sum of the individual resistances


"R=R_1+R_2+R_3"

"I=\dfrac{E_t}{R_1+R_2+R_3}"

The voltage drop in "R_1" is


"U_1=IR_1=\dfrac{R_1}{R_1+R_2+R_3}E_t"

"U_1=\dfrac{12\ Ohm}{12\ Ohm+20\ Ohm+5\ Ohm}\cdot24 \ V"

"=\dfrac{288}{37}\ V\approx7.784\ V"